OpenStudy (anonymous):

find any x intercepts find any veritcal asymptote(s) find any y intercepts find any horizontal or oblique asymptotes determine wherther or not the grpah cosses the horizontal or oblique asymptot. if so determine where F(x)=6x^3+3x^2-3x over 2x^3-7x^2+7x-2

OpenStudy (dumbcow):

a 3x factors on top 3x(2x^2+x-1)=0 3x(2x-1)(x+1)=0 therefore x-intercepts at -1,0,1/2

OpenStudy (anonymous):

okay what else?

OpenStudy (dumbcow):

same degree on top and bottom, so horizontal asymptote at y = 6/2 = 3

OpenStudy (dumbcow):

for vertical asymptotes, you have to find zeroes of the cubic function on bottom

OpenStudy (dumbcow):

x=1 is a solution, factor out a (x-1) using synthetic division 1 | 2 -7 7 -2 2 -5 2 2 -5 2 0 2x^2 -5x +2 =0 (2x-1)(x-2) = 0 vertical asymptotes at 1/2,1,2

OpenStudy (dumbcow):

actually there is a point of discontinuity at x=1/2 the limit of f(x) at x=1/2 is 3